A rectangle has a length 2 meters less than twice its width. When 5 meters is added to the width, the result is a square with an area of 144 square meters. Find the dimensions of the original rectangle

Answer by checkley77(12844)

L=2W-2
(2W-2)(W+5)=144
2W^2-2W+10W-10=144
2W^2+8W-10-144=0
2W^2+8W-154=0
2(W^2+4W-77)=0
2(W+11)(W-7)=0
W-7=0
W=7 ANS FOR THE ORIGINAL WIDTH.
7+7=12 IS THE EXPANDED WIDTH.
L=2*7-2
L=14-2=12 ANS. FOR THE LENGTH.
PROOF:
12*12=144
144=144